OpenStudy (anonymous):

Q) e^2y + xy = 4x + 1. find dy/dx and equation of the tangent to the curve.

OpenStudy (isaiah.feynman):

At what point?

OpenStudy (anonymous):

the origin.

OpenStudy (isaiah.feynman):

(0,0). Differentiate the function implicitly. Then evaluate it at (0,0).

OpenStudy (anonymous):

Can anyone help with differentiating the function?

OpenStudy (isaiah.feynman):

Not quite.

OpenStudy (anonymous):

Ok my latest attempt I get: dy/dx = 4 - y / e^(2y) + x I hope that's correct?

OpenStudy (isaiah.feynman):

Yes, but you are missing a 2 in front of e.

OpenStudy (anonymous):

Yep - sorry typo. Finally - hooray! Now for the next part of the Qu.

OpenStudy (isaiah.feynman):

Simply substitute, 0 for x and y and see the outcome. Well the outcome will be the slope.

OpenStudy (isaiah.feynman):

Of that line.

OpenStudy (anonymous):

That gives me 4/3.

OpenStudy (isaiah.feynman):

Where would 3 come from??

OpenStudy (anonymous):

oops sorry, 4/2 = 2...

OpenStudy (isaiah.feynman):

That's the SLOPE of that line. Remember how to find the equation of a line when you have the slope and point? From algebra classes in high school.

OpenStudy (anonymous):

Yep the equation part is fine. Thanks.

OpenStudy (isaiah.feynman):

Good good.

OpenStudy (anonymous):

y-0=2(x-0) y=2x

OpenStudy (isaiah.feynman):

Nice.

OpenStudy (anonymous):

Thanks.

OpenStudy (isaiah.feynman):

Welcome.